OpenStudy (anonymous):

solve the substitution method 8m+n=35 m-8n=45 answer needs to be in ordered pair

OpenStudy (blacksteel):

Okay, lets rewrite the second equation as m = 45 + 8n. Then we substitute this into the first equation to get 8(45 + 8n) + n = 35, or 360 + 65n = 35, or 65n = -325, or n = -5. Then we can use n = -5 to solve the second equation: m - 8*(-5) = 45, or m + 40 = 45, or m = 5.

OpenStudy (blacksteel):

In ordered pair form, your answer is (m,n) = (5,-5).

OpenStudy (anonymous):

frm first eqn u get n = 35 -8m putting this in second eqn u get m -8(35-8m) = 45 m - 280 + 64m = 45 65m = 45 + 280 m = 325/65 m = 5 substitute m = m in first eqn 8(5) + n = 35 40 + n = 35 n = 35 - 40 n = -5 so m = 5 and n = -5